Use two equations in two variables to solve the application.
A boat can travel 24 miles downstream in 2 hours and can make the return trip in 4 hours. Find the speed of the boat in still water.

Answer by ikleyn(52787)

.
Let x be the boat speed in still water (in miles per hoyr).
Let y be the speed of the current.


Then from the condition


x + y = 24/2 = 12,    (1)
x - y = 24/4 =  6     (2)


To solve the system, add the equations. You will get

2x = 12 + 6 = 18  ====>  x = 18/2 = 9.


Answer.   The boat speed in still water is 9 mph.

Solved.